Vulcan Company’s contribution format income statement for June is as follows:

Vulcan Company
Income Statement
For the Month Ended June 30
Sales $ 900,000
Variable expenses 408,000
Contribution margin 492,000
Fixed expenses 465,000
Net operating income $ 27,000

Management is disappointed with the company’s performance and is wondering what can be done to improve profits. By examining sales and cost records, you have determined the following:

The company is divided into two sales territories—Northern and Southern. The Northern Territory recorded $400,000 in sales and $168,000 in variable expenses during June; the remaining sales and variable expenses were recorded in the Southern Territory. Fixed expenses of $212,000 and $130,000 are traceable to the Northern and Southern Territories, respectively. The rest of the fixed expenses are common to the two territories.

The company is the exclusive distributor for two products—Paks and Tibs. Sales of Paks and Tibs totaled $200,000 and $200,000, respectively, in the Northern territory during June. Variable expenses are 25% of the selling price for Paks and 59% for Tibs. Cost records show that $90,000 of the Northern Territory’s fixed expenses are traceable to Paks and $40,000 to Tibs, with the remainder common to the two products.

1-a. Prepare contribution format segmented income statements for the total company broken down between sales territories.

1-b. Prepare contribution format segmented income statements for the Northern Territory broken down by product line.

Prepare contribution format segmented income statements for the total company broken down between sales territories. (Round the percentage answers to one decimal place (i.e .1234 should be entered as 12.3).)

Sales Territory
Total Company Northern Southern
Amount % Amount % Amount %
Sales      
Variable expenses
Contribution margin
Traceable fixed expenses
Territorial segment margin   
Common fixed expenses
Net operating income

Prepare contribution format segmented income statements for the Northern Territory broken down by product line. (Round the percentage answers to one decimal place (i.e .1234 should be entered as 12.3).)

Product Line
Northern Territory Paks Tibs
Amount % Amount % Amount %
Sales
Variable expenses
Contribution margin
Traceable fixed expenses
Product line segment margin   
Common fixed expenses
Sales territory segment margin
Answer #1

Sales for the southern territory = 900,000 - 400,000 = $500,000

Common fixed expenses = 465,000 - 342,000 = $123,000

Sales Territory Total Company Northern Southern Amount % Amount Amount Sales 900,000 400,000 | 44.4 500,000 55.5 58.8 Variabl

Common fixed expenses = 342,000 - 130,000 = $212,000

TOTAL Product Line Northern Territory Paks Tibs Amount % Amount % Amount % Sales 400.000 200,000 50 200,000 50 Variable expen